OpenStudy (anonymous):

Solve x − 2y = 6 3x − 6y = 18 (2, −2) (3, negative three halves) No Solutions Infinitely Many Solutions

OpenStudy (arabpride):

turn into y = mx + b form x − 2y = 6

OpenStudy (luigi0210):

Solve by sub \(\large x=2y+6\) \(\large 3x-6y=18\) \[\LARGE 3(2y+6)-6y=18\] Distribute and solve.

OpenStudy (arabpride):

then work from there

OpenStudy (solomonzelman):

divide the 3rd equation (each thing in it) by 3. it will turn into the 1st equation.

OpenStudy (luigi0210):

._.

OpenStudy (arabpride):

x - 2y = 6 - x -x -2y = -x +6 /-2 /-2 >>>>>>y = x/2 - 3<<<<<<

OpenStudy (solomonzelman):

What ?

OpenStudy (whpalmer4):

Yes, that's correct — the solution is the equation of a line. What does that mean about the number of solutions?

OpenStudy (luigi0210):

If you solve it by sub you get.. \(\large 18=18\) Do you know what that means? >_>

OpenStudy (arabpride):

y = mx + b form: 3x − 6y = 18 -3x -3x -6y = -3x + 18 /-6 /-6 >>>>>>y = 1/2x - 3<<<<< thats our second equation^^

OpenStudy (arabpride):

y = x/2 - 3 y = 1/2x - 3 Since they are on the same line ~ they can go on & on so: Infinitely Many Solutions

OpenStudy (arabpride):

@AyyoDutches

OpenStudy (luigi0210):

btw.. \(\LARGE \frac{x}{2} \neq \frac{1}{2x}\)

OpenStudy (whpalmer4):

yeah, that's sloppy writing, should be y = (1/2)x-3

OpenStudy (solomonzelman):

luigi, if x=1, then yes ;)

OpenStudy (luigi0210):

Why couldn't you guys just use substitution? ლ(ಠ_ಠლ) And I see what you did there Solomon xD

OpenStudy (solomonzelman):

(*_*)

OpenStudy (anonymous):

>Confused<

OpenStudy (whpalmer4):

@AyyoDutches You have two equations which turn out to represent the same line. As any point that is on both lines is a solution, that means that any point on either line is a solution, and there are infinitely many points on a line, so there are infinitely many solutions.

OpenStudy (arabpride):

~ the answer is Infinitely Many Solutions ~ because of the explanations above XD

OpenStudy (whpalmer4):

If you had two parallel lines, you would have no solutions. An example would be y = x+5 y = x+6 same slope, so parallel, but no point satisfies both equations, so there are no solutions

OpenStudy (whpalmer4):

most often when solving such problems, the lines intersect (once only, if straight lines), and thus the point of intersection is the only point that satisfies all of the equations in the system, giving you just one solution